# Ego-to-Exo: Interfacing Third Person Visuals from Egocentric Views in Real-time for Improved ROV Teleoperation

### v1

@@ Abstract
Underwater ROVs (Remotely Operated Vehicles) are unmanned submersible vehicles designed for exploring and operating in the depths of the ocean. Despite using high-end cameras, typical teleoperation engines based on first-person (egocentric) views limit a surface operator’s ability to maneuver and navigate the ROV in complex deep-water missions. In this paper, we present an interactive teleoperation interface that (i) offers on-demand “third”-person (exocentric) visuals from past egocentric views, and (ii) facilitates enhanced peripheral information with augmented ROV pose in real-time. We achieve this by integrating a 3D geometry-based Ego-to-Exo view synthesis algorithm into a monocular SLAM system for accurate trajectory estimation. The proposed closed-form solution only uses past egocentric views from the ROV and a SLAM backbone for pose estimation, which makes it portable to existing ROV platforms. Unlike data-driven solutions, it is invariant to applications and waterbody-specific scenes. We validate the geometric accuracy of the proposed framework through extensive experiments of 2-DOF indoor navigation and 6-DOF underwater cave exploration in challenging low-light conditions. We demonstrate the benefits of dynamic Ego-to-Exo view generation and real-time pose rendering for remote ROV teleoperation by following navigation guides such as cavelines inside underwater caves. This new way of interactive ROV teleoperation opens up promising opportunities for future research in underwater telerobotics.
@@ I Introduction
Unmanned submersible vehicles such as ROVs (Remotely Operated Vehicles) play a crucial role in subsea inspection, remote surveillance, and underwater cave exploration [1, 2, 3]. They are particularly useful in inspecting deep-water structures and surveying confined spaces that are beyond the reach of human scuba divers [4, 5]. In a typical mission, ROVs are controlled by human operators from a surface vessel, who are responsible for the safe and efficient maneuvering [6, 7]. The control consoles for teleoperation typically offer real-time data such as the egocentric video feed, pose, velocity, depth, etc. State-of-the-art ROVs can also include autonomous features for atomic tasks such as hovering [8], following navigation guidelines (e.g., cavelines inside underwater caves [9, 10]), gripping objects [11, 12], trajectory estimation for visual servoing, etc.
@@
Figure 1: The proposed Ego-to-Exo teleoperation interface is demonstrated for an underwater cave exploration scenario with an ROV. The traditional console interfaces are based on egocentric views (top left), which is limiting and disorienting to a surface operator in noisy low-light conditions. Our method generates on-demand exocentric views from a given EOB (Eye On the Back) viewpoint, i.e., third-person views from behind the ROV. It also projects the ROV point clouds to augment a comprehensive exocentric perspective with real-time ROV pose. As shown, this interactive EOB viewpoint selection option is integrated into a standard BlueROV2 console for a significantly improved teleoperation experience.
In this paper, we address these issues by introducing an AR (augmented reality) inspired ROV teleoperation interface that generates third-person (exocentric) perspectives as well as provides interactive control choices for viewpoint selection. As shown in Fig. 1, the proposed console can generate multiple exocentric views from past egocentric images, with the ROV’s point cloud projected on the images as if it were taken by a third person immediately behind the robot. Our earlier work introduced the idea of EOB (Eye On the Back) visuals [14], which produces a single third-person view from immediately behind the ROV to facilitate better teleoperation. This work develops the AR-based end-to-end pipeline for generating on-demand exocentric perspectives given any EOB viewpoints. Additionally, we integrate the feature for geometrically accurate ROV positioning into those views for an interactive ROV teleoperation experience in challenging underwater applications. With these features, an operator can slide across multiple EOB viewpoints for the best view, see where the ROV is, and get comprehensive peripheral information for safe and efficient teleoperation.
@@
The significance of this work resides in the simplicity of a 3D-geometric formulation of the Ego-to-Exo problem, and its integration into a real-time SLAM system. A monocular vision-only pipeline is chosen to ensure generalized utility and computational efficiency. The proposed Ego-to-Exo solution, EOB viewpoint-based parameterization, and ROV point cloud projection – are carried out by closed-form solutions to ensure real-time performance. As opposed to data-driven approaches for exocentric view synthesis, the proposed framework is invariant to the changes in waterbody style, scene geometry, and application scenario – making it transferable to existing underwater ROV platforms.
We demonstrate the performance and utility of the proposed Ego-to-Exo framework through a series of experiments. First, we prove the geometric validity of our estimation with 2D indoor navigation scenarios utilizing a TurtleBot4. We quantify the geometric validity based on ground plane estimation errors and reprojection errors of known reference points in the scene. Next, we conduct experiments for autonomous underwater cave exploration scenarios where a BlueROV2 is teleoperated inside underwater caves and grotto systems at various geographical locations. Remote teleoperation in such cave systems includes unique challenges such as low visibility, turbid water conditions, moving shadow effects, and hazy blind spots due to artificial lights. In these scenarios, we demonstrate how the generated Ego-to-Exo views, the augmented robot pose, and the updated map can facilitate improved underwater ROV teleoperation.
In particular, we discuss the results and observations for a variety of challenging cases of caveline detection and following in noisy low-light conditions inside multiple underwater caves and grotto systems. We demonstrate that the generated exocentric views embed significantly more information and global context about the scene for safe and efficient robot control. The operators can adjust viewpoints to visualize augmented ROV poses on the map in real-time – facilitating an interactive teleoperation experience.
@@ II-A ROV Deployment & Teleoperation
Subsea Structure Inspection
Underwater structures operate in challenging remote environments and are susceptible to various hazards with significant economic and environmental consequences. Corrosion/erosion combined with tensile stresses caused by waves and seismic activities lead to cracking, leakage, and malfunction which are regularly inspected by ROVs [1, 17]. ROVs are also deployed for 3D mapping and acoustic profiling of marine structures [18, 19].
Underwater infrastructure maintenance & security
The underwater energy sector [20] and submerged data centers [21, 22] are experiencing rapid growth with increasing demand for regular maintenance and surveillance operations. Once ROVs are teleoperated to the site, they need to perform various autonomous/semi-autonomous tasks to inspect, monitor, and ensure safety and integrity of those infrastructures.
Marine archaeology and cave exploration
Exploring underwater heritage sites [23] and caves [4] are important to unravel archaeological history as well as for water resource management. Intelligent ROV capabilities enable effective mapping, sampling, and exploration of remote sites and challenging overhead environments. Moreover, seasonal monitoring of oceanographic features is performed for photometric studies as well [24, 25].
@@ II-B Third-Person Views for ROV Teleoperation
A common issue reported by ROV operators is that using a remote vision platform for teleoperation is like looking through a “soda straw” [26]. This is because the typical ROV controller interfaces are based on egocentric first person camera views – which provide no peripheral vision, resulting in significantly reduced situational awareness [27, 28]. As such, there have been multiple methods proposed to mitigate the issue and provide the operator with more visual data [29].
Two types of commonly used approaches exist for generating an exocentric view for unmanned ground and aerial vehicles. The first method uses an external camera to provide the exocentric view such as using a fixed camera to record the vehicle’s trajectory from a distance [30], mounting the camera on a UAV (unmanned aerial vehicle) that follows above the target [31, 32, 33], attaching the camera to an elevated position on the robot [34], utilizing a follower ROV with a camera to provide the external view [35], or using a fish-eye camera to create a top-down perspective [36]. The second method utilizes integrated sensors on the ROV, such as LiDAR (Light Detection and Ranging) to generate a point cloud of the surrounding environment [37, 38] and use that toward creating an augmented/virtual environment for real-time interfacing and teleoperation [39, 40, 41].
Adapting the aforementioned methods from terrestrial or aerial domains to underwater environments presents inherent challenges. First, sending diver-robot teams [42] is not always an option in complex deep-water missions – which are the major use cases for ROVs. Secondly, installing an external visual system requires significant hardware modifications, e.g., they need to be rugged and pressure-sealed with the rest of the robot, re-calibrated for buoyancy and motion dynamics, and additional tether integration for high-speed exocentric data transfer. Even with all the structural modifications, an external camera will provide only one additional third-person perspective. We attempt to address these issues by synthesizing geometrically accurate exocentric views into a monocular SLAM pipeline given a dynamic viewpoint.
@@ III-B Generating Exocentric Perspective
As shown in Fig. 2(a), we use the ROV point cloud set $\mathbf{P}$ of size $3\times m$ as prior. These $m$ points are transformed from current camera pose ${}^{w}_{c}\mathbf{T}$ to reference camera pose ${}^{w}_{r}\mathbf{T}$ using:
@@ IV-B Proof of Concept: 2D Indoor Navigation
Geometric validation: reprojection error analysis. We first evaluate reprojection errors of known reference points from the generated Ego-to-Exo views and estimated ROV pose. As shown Fig. 3, we use standard checkerboard corners as reference points from egocentric views and then evaluate the reprojection errors for those points from exocentric views. This test is iterated over different sets of past egocentric images, each corresponding to a different EOB distance. As shown in Fig. 3(a), a checkerboard is viewed from different EOB distances (further back into the past) indicated by the parameter $f$. More specifically, $f$ is the number of frames between the current egocentric view and the selected EOB view. The corresponding reprojection error is plotted in Fig. 3(b), which shows how the estimation is accurate for lower values of $f$, and gradually degenerates for $f>100$. This is consistent with our visual observation of the projected ROV point cloud, i.e., it is on the ground plane with accurate orientation based on the SLAM trajectory estimates.
@@
Geometric validation: ground plane estimation. We also visualize the accuracy of the estimated ground plane in these experiments for qualitative assessments. Fig. 3(b) illustrates two cases: one for $f=70$ with a low reprojection error, and another for $f=260$ with a high error. As seen, the estimated ground plane (and the drawn sample cube) validates the geometric accuracy for $f=70$ case. On the other hand, a misaligned ground plane for $f=260$ case demonstrates the underlying error in pose estimation as well as in the reprojection process. Essentially, the geometric accuracy of our proposed Ego-to-Exo algorithm depends on the trajectory estimation performance of the SLAM system. More importantly, a teleoperator can make the right viewpoint selection in real-time for optimal Ego-to-exo rendering.
@@
Computational efficiency. To ensure real-time execution in resource-constrained edge devices onboard standard ROV platforms, we analyze the computational complexity of the Ego-to-Exo algorithm for different configurations. Table I shows the memory requirement of our algorithm for different choices of buffer size. The memory footprint is less than $300$ MB for a buffer size of up to $180$ frames, making it highly efficient. Table II demonstrates that the end-to-end framework maintains a consistent output rate of over $25$ FPS (frames per second), making it suitable for integration in existing teleoperation engines. In fact, the Ego-to-Exo algorithm only adds $6.8\%$ overhead on the SLAM backbone for the global map update.
@@ IV-C Field Deployment: 3D Underwater Cave Exploration
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV possesses full 6-DOF motion. While the $roll$ motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as a guide [46]. The mission objective is to navigate the ROV $75$-$300$ feet deep inside the cave through its complex structures, and then safely return it to the surface. In addition to evaluating the geometric accuracy and robustness, we consider how informative the generated Ego-to-Exo views are compared to traditional consoles for underwater ROV teleoperation.
Real-time map update and teleoperation. In addition to the Ego-to-Exo view generation and ROV pose rendering, our framework simultaneously updates a 3D map with extracted feature points from the SLAM system. Fig. 4 shows an ROV’s trajectory mapped during an underwater cave mission in Devil’s Springs, Florida. As seen, the generated Ego-to-Exo views embed significantly more peripheral information about the scene. The exocentric view of the ROV pose and its relative distance from cave walls or overhead obstacles are useful to surface operators for obstacle avoidance and efficient decision-making. Additionally, the 3D map shows the ROV’s past trajectory and its current pose which are useful to analyze the mission progress, which is not possible in traditional teleoperation consoles. Such a global view of the trajectory map is also useful during emergency evacuation and recovery. Beyond underwater cave exploration, these features will be crucial in ROV-based subsea surveillance and search-and-rescue operations as well.
Accuracy and robustness. Due to the complex scene geometry and absence of a smooth ground plane inside underwater caves, we adopt a homography estimation approach for performance validation. As shown in Fig. 5, April-Tag [47] corners are used as reference points for reprojection. Specifically, we compute the homography transformation between the egocentric and synthesized exocentric views to visualize the reprojection errors. As Fig. 5 shows, we use a sample 2D logo and project it onto the reference April-Tag surface using the homographic transform. The unskewed planar projection validates the accuracy of the Ego-to-Exo pose estimation and point cloud rendering processes.

### v2

@@ Abstract
Underwater ROVs (Remotely Operated Vehicles) are unmanned submersible vehicles designed for exploring and operating in the depths of the ocean. Despite using high-end cameras, typical teleoperation engines based on first-person (egocentric) views limit a surface operator’s ability to maneuver the ROV in complex deep-water missions. In this paper, we present an interactive teleoperation interface that enhances the operational capabilities via increased situational awareness. This is accomplished by (i) offering on-demand “third”-person (exocentric) visuals from past egocentric views, and (ii) facilitating enhanced peripheral information with augmented ROV pose information in real-time. We achieve this by integrating a 3D geometry-based Ego-to-Exo view synthesis algorithm into a monocular SLAM system for accurate trajectory estimation. The proposed closed-form solution only uses past egocentric views from the ROV and a SLAM backbone for pose estimation, which makes it portable to existing ROV platforms. Unlike data-driven solutions, it is invariant to applications and waterbody-specific scenes. We validate the geometric accuracy of the proposed framework through extensive experiments of 2-DOF indoor navigation and 6-DOF underwater cave exploration in challenging low-light conditions. A subjective evaluation on $15$ human teleoperators further confirms the effectiveness of the integrated features for improved teleoperation. We demonstrate the benefits of Ego-to-Exo view generation and real-time pose rendering for remote ROV teleoperation by following navigation guides such as cavelines inside underwater caves. This new way of interactive ROV teleoperation opens up promising opportunities for future research in subsea telerobotics.
@@ I Introduction
Unmanned submersible vehicles such as ROVs (Remotely Operated Vehicles) play a crucial role in subsea inspection, remote surveillance, and underwater cave exploration [1, 2, 3]. They are particularly useful in inspecting deep-water structures and surveying confined spaces that are beyond the reach of human scuba divers [4, 5]. In a typical mission, ROVs are controlled by human operators from a surface vessel, who are responsible for the safe and efficient maneuvering of the vehicle [6, 7]. The control consoles for teleoperation typically offer real-time data such as the egocentric video feed, pose, velocity, depth, etc. State-of-the-art ROVs can also include autonomous features for atomic tasks such as hovering [8], following navigation guidelines (e.g., cavelines inside underwater caves [9, 10]), gripping objects [11, 12], trajectory estimation, etc.
@@
Figure 1: The proposed Ego-to-Exo teleoperation interface is demonstrated for an underwater cave exploration scenario with an ROV. The traditional console interfaces are based on egocentric views (top left), which is limiting and disorienting to a surface operator in noisy low-light conditions. Our method generates on-demand exocentric views from a given EOB (Eye On the Back) viewpoint, i.e., third-person views from behind the ROV. This interactive EOB viewpoint-selection option is integrated into a standard BlueROV2 console for a significantly improved teleoperation experience.
In this paper, we address these issues by introducing an AR (augmented reality) inspired ROV teleoperation interface that generates third-person (exocentric) perspectives as well as provides interactive control choices for viewpoint selection. As shown in Fig. 1, the proposed console can generate multiple exocentric views from past egocentric images, with a virtual ROV projected on the images as if it were taken by a third person following the robot. Our earlier work introduced the idea of EOB (Eye On the Back) visuals [14], which produces a single third-person view from immediately behind the ROV to facilitate better teleoperation. This work develops the AR-based end-to-end pipeline for generating on-demand exocentric perspectives given any EOB viewpoints. Additionally, we integrate the feature for geometrically accurate ROV positioning into those views for an interactive ROV teleoperation experience in challenging underwater applications. With these features, an operator can slide across multiple EOB viewpoints for the best view, see where the ROV is, and get comprehensive peripheral information for safe and efficient teleoperation.
@@
The significance of this work resides in the simplicity of a 3D-geometric formulation of the Ego-to-Exo problem, and its integration into a real-time SLAM system. A monocular vision-only pipeline is chosen to ensure generalized utility and computational efficiency. The proposed Ego-to-Exo solution, EOB viewpoint-based parameterization, and ROV point cloud projection – are carried out by closed-form solutions to ensure real-time performance. As opposed to data-driven approaches for exocentric view synthesis, the proposed framework is invariant to the changes in waterbody style, scene geometry, and application scenario – making it transferable to any underwater ROV platforms.
We demonstrate the performance and utility of the proposed Ego-to-Exo framework through a series of experiments. First, we prove the geometric validity of our estimation with 2D indoor navigation scenarios utilizing a TurtleBot4. We quantify the geometric validity based on ground plane estimation errors and reprojection errors of known reference points in the scene. Next, we conduct experiments for underwater cave exploration scenarios where a BlueROV2 is teleoperated inside underwater caves and grotto systems at various geographical locations. Remote teleoperation in such cave systems includes unique challenges such as low visibility, turbid water conditions, moving shadow effects, and hazy blind spots due to artificial lights. We evaluate the usefulness of the proposed TeleOp console with $15$ human subjects in such challenging scenarios. They use the interactive features such as viewpoint selection to visualize augmented ROV poses on the map in real-time. The subjects rank the complexity, efficiency, and other aspects of the system in a questionnaire as well as provide open-ended feedback regarding their cognitive workload.
We further discuss the observations for a variety of challenging cases of caveline detection and following in noisy low-light conditions inside multiple underwater caves. We demonstrate that the generated exocentric views, the augmented robot pose, and the rendered map embed significantly more information and global context about the scene to facilitate safe and efficient robot control.
@@ II-A ROV Deployment & Teleoperation
Subsea structure inspection. Underwater structures operate in challenging remote environments and are susceptible to various hazards with significant economic and environmental consequences. Corrosion/erosion combined with tensile stresses caused by waves and seismic activities lead to cracking, leakage, and malfunction which are regularly inspected by ROVs [1, 17]. ROVs are also deployed for 3D mapping and acoustic profiling of marine structures [18, 19].
Underwater infrastructure maintenance & security. The underwater energy sector [20] and submerged data centers [21, 22] are experiencing rapid growth with increasing demand for regular maintenance and surveillance operations. Once ROVs are teleoperated to the site, they need to perform various autonomous and semi-autonomous tasks to inspect, monitor, and ensure safety and integrity of the infrastructures.
Marine archaeology and cave exploration. Exploring underwater heritage sites [23] and caves [4] are important to unravel archaeological history as well as for water resource management. Intelligent ROV capabilities enable effective mapping, sampling, and exploration of remote sites and challenging overhead environments. Moreover, seasonal monitoring of oceanographic features is performed for photometric studies as well [24, 25, 26].
@@ II-B Third-Person Views for ROV Teleoperation
A common issue reported by ROV operators is that using a remote vision platform for teleoperation is like looking through a “soda straw” [27]. This is because the typical ROV controller interfaces are based on egocentric first person camera views – which provide no peripheral vision, resulting in significantly reduced situational awareness [28, 29]. As such, there have been multiple methods proposed to mitigate the issue and provide the operator with more visual data.
Two types of commonly used approaches exist for generating an exocentric view for unmanned ground and aerial vehicles. The first method uses an external camera to provide the exocentric view such as using a fixed camera to record the vehicle’s trajectory from a distance [30], mounting the camera on a UAV (unmanned aerial vehicle) that flies above the target [31, 32, 33, 34], attaching the camera to an elevated position on the robot [35], utilizing a follower ROV with a camera to provide the external view [36], or using a fish-eye camera to create a top-down perspective [37]. The second method utilizes integrated sensors on the robot, such as LiDAR (Light Detection and Ranging) to generate a point cloud of the surrounding environment [38, 39] and use that toward creating an augmented/virtual environment for interfacing and teleoperation [40, 41, 42].
Adapting the aforementioned methods from terrestrial or aerial domains to underwater environments presents inherent challenges. Firstly, sending diver-robot teams [43] is not always an option in complex deep-water missions – which are the majority of use cases for ROVs. Secondly, UGVs that utilize past egocentric views [44, 45] primarily rely on GPS-based localization that does not apply to GPS-denied underwater environment. Unlike underwater ROVs, ground vehicles generally operate on a 2D plane with limited pitch and roll variations over rough terrain. Thirdly, installing an external visual system requires significant hardware modifications, e.g., they need to be rugged and pressure-sealed, re-calibrated for buoyancy and motion dynamics, and additional tether integration for high-speed exocentric data transfer. Even with all the structural modifications, an external camera will provide a single additional third-person perspective. We attempt to address these issues by synthesizing geometrically accurate exocentric views into a monocular SLAM pipeline given a dynamic viewpoint.
@@ III-B Generating Exocentric Perspective
We use the ROV point cloud model $\mathbf{P}$ of size $3\times m$ as prior. These $m$ points are transformed from current camera pose ${}^{w}_{c}\mathbf{T}$ to reference camera pose ${}^{w}_{r}\mathbf{T}$ using:
@@ IV-B Proof of Concept: 2D Indoor Navigation
Geometric validation: reprojection error analysis. We first evaluate the reprojection errors of known reference points from the generated Ego-to-Exo views and the estimated ROV pose. As shown Fig. 3, we use standard checkerboard corners as reference points from egocentric views and then evaluate the reprojection errors for those points from exocentric views. This test is iterated over different sets of past egocentric images, each corresponding to a different EOB distance. As shown in Fig. 3(a), a checkerboard is viewed from different EOB distances (further back into the past) indicated by the parameter $f$. More specifically, $f$ is the number of frames between the current egocentric view and the selected EOB view. The corresponding reprojection error is plotted in Fig. 3(b), which shows how the estimation is accurate for lower values of $f$, and gradually degenerates for $f>100$. This is consistent with our visual observation of the projected ROV point cloud, i.e., it is on the ground plane with accurate orientation based on the SLAM trajectory estimates.
@@
Geometric validation: ground plane estimation. We also visualize the accuracy of the estimated ground plane in these experiments for qualitative assessments. Fig. 3(b) illustrates two cases: one for $f=70$ with a low reprojection error, and another for $f=260$ with a high error. As seen, the estimated ground plane (and the drawn sample cube) validates the geometric accuracy for the $f=70$ case. On the other hand, a misaligned ground plane for the $f=260$ case demonstrates the underlying error in pose estimation as well as in the reprojection process. Essentially, the geometric accuracy of our proposed Ego-to-Exo algorithm depends on the trajectory estimation performance of the SLAM system. More importantly, a teleoperator can make the right viewpoint selection in real-time for optimal Ego-to-exo rendering.
@@
Computational efficiency. We analyze the computational complexity of the Ego-to-Exo algorithm for different configurations to ensure real-time execution in resource-constrained edge devices onboard standard ROV platforms. Table I shows the memory requirement of our algorithm for different choices of buffer size. The memory footprint is less than $300$ MB for a buffer size of up to $180$ frames, making it highly efficient. Table II demonstrates that the end-to-end framework maintains a consistent output rate of over $25$ FPS (frames per second), making it suitable for integration in existing teleoperation engines. In fact, our algorithm only adds $6.8\%$ overhead on the SLAM backbone for the global map update.
@@ IV-C Field Deployment: 3D Underwater Cave Exploration
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV performs full 6-DOF motions. While the $roll$ motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as a guide [49]. The mission objective is to navigate the ROV $75$-$300$ feet deep inside the cave through its complex structures, and then safely return it to the surface. In addition to evaluating the geometric accuracy and robustness, we consider how informative the generated Ego-to-Exo views are compared to traditional consoles for ROV teleoperation.
Real-time map update and teleoperation. In addition to the Ego-to-Exo view generation and ROV pose rendering, our framework simultaneously updates a 3D map with extracted feature points from the SLAM system. Fig. 4 shows an ROV’s trajectory mapped during an underwater cave mission in Devil’s Springs, Florida. As seen, the generated Ego-to-Exo views embed significantly more peripheral information about the scene. The exocentric view of the ROV pose and its relative distance from cave walls or overhead obstacles are useful to surface operators for obstacle avoidance and efficient decision-making. Additionally, the 3D map shows the ROV’s past trajectory and its current pose which are useful to analyze the mission progress, which is not possible in traditional teleoperation consoles. Such a global view of the trajectory is also useful during emergency evacuation and recovery. Beyond cave exploration, these features will be crucial in ROV-based subsea surveillance and search-and-rescue operations as well.
@@
Accuracy and robustness. Due to the complex scene geometry and absence of a smooth ground plane inside underwater caves, we adopt a homography estimation approach for the performance validation. As shown in Fig. 5, April-Tag [50] corners are used as reference points for reprojection. Specifically, we compute the homography transformation between the egocentric and synthesized exocentric views to visualize the reprojection errors. We use a sample 2D logo and project it onto the reference April-Tag surface using the homographic transform. The unskewed planar projection validates the accuracy of the Ego-to-Exo pose estimation and point cloud rendering processes.
TABLE III: $15$ subjects provide their feedback to the following two sets of questions: Question $1-10$ from SUS [51], and the rest are custom designed. Response to each question is scaled from $1$ (strongly disagree) to $5$ (strongly agree).
# Question mean, std
System Usability Scale (SUS)[51]
1 I think that I would like to use this system frequently. 4.3,absent,~{},0.6
2 I found the system unnecessarily complex. 2.0,absent,~{},0.7
3 I thought the system was easy to use. 4.3,absent,~{},0.4
4 I think that I would need the support of a technical person to be able to use this system. 2.0,absent,~{},0.6
5 I found the various functions in this system were well integrated. 4.0,absent,~{},0.8
6 I thought there was too much inconsistency in this system. 2.3,absent,~{},0.7
7 I would imagine that most people would learn to use this system very quickly. 4.4,absent,~{},0.5
8 I found the system very cumbersome to use. 2.0,absent,~{},0.6
9 I felt very confident using the system. 3.7,absent,~{},0.7
10 I needed to learn a lot of things before I could get going with this system. 1.4,absent,~{},0.5
Custom Questions
11 The proposed exocentric view is beneficial for ROV teleoperation. 4.5,absent,~{},0.5
12 I found the EOB distance tuning feature useful to get the best view. 4.5,absent,~{},0.5
13 The generated 3D map provides a better understanding of the ROV’s global location and its surroundings. 4.6,absent,~{},0.9
@@
IV-D Subjective User Study
The user study is conducted with multiple underwater cave exploration data. A BlueROV2 recorded egocentric video feeds inside the caves at up to $100$-meter penetrations. Later on, $15$ participants evaluate the ease of operation with our developed console and compare it to traditional consoles. Their feedback is recorded using the System Usability Scale (SUS) [51], with our interface achieving an average SUS score of $77.5$. We also formulate an independent set of questions that reflect the teleoperator’s preference for the novel features of our method. The individual questions and corresponding scores are presented in Table III. Some key observations from this study are listed below.
  1. 1.
    The obtained SUS score is fairly above median (score: $68$) and is considered Good for user experience; it is slightly below the Excellent (score: $80.3$) category.
  2. 2.
    Post-operation feedback from our ROV operators suggests that the exocentric views are more useful for safe ROV maneuvers.
  3. 3.
    The synthesized 3D map provides a better sense of the ROV’s global location and improves spatial awareness of the teleoperators.
  4. 4.
    The operators report a significantly less workload (perceived cognitive load) in conducting complex tasks such as object following and structure mapping.